# Predicting Time Pressure of Powered Two-Wheeler Riders for Proactive Safety Interventions

### v1

@@ VII-B Application in ITS Interventions
The predicted TP probabilities are thresholded to classify riders into discrete states (NTP, LTP, HTP), as shown in Table XII and Fig. 8. This state-based risk index reframes ITS from passive monitoring to proactive, preventive interventions. Under the Safe System Approach (SSA), the emphasis is not only on preventing crashes but also on minimizing harm when crashes occur. By integrating MTPS-predicted TP states into adaptive human–machine interfaces (haptics, audio, AR) and vehicle-to-infrastructure (V2I) signaling (e.g., road side unit-based speed guidance, dynamic signal phasing, or targeted alerts), interventions can be triggered before risky behavior escalates. This enables systems to modulate kinetic energy exposure in real time (e.g., slowing riders earlier, guiding safer gap acceptance, or synchronizing vehicle–infrastructure interactions) thereby reducing the severity of crashes even when they are unavoidable. MTPS thus translates cognitive stress estimation into actionable intelligence for SSA-compliant ITS, where thresholds are aligned with both crash likelihood and energy transfer risk, supporting holistic harm reduction strategies.
@@ VIII Discussions and Conclusions
Building on these empirical insights that demonstrate the systemic impact of TP on riding safety, we developed MTPS, a novel DL architecture that achieves 91.53% accuracy and 98.93% ROC AUC in classifying TP states, outperforming eight diverse baselines. Beyond classification accuracy, MTPS demonstrates significant practical utility by improving downstream crash-risk prediction, integrating MTPS-predicted TP into an Informer-based model increases accuracy from 91.25% to 93.51%, approaching oracle-level performance (93.72%). Additional validation confirms perfect HTP discrimination (AUC = 1.00) and reliable confidence calibration essential for real-world deployment.
In conclusion, MTPS is a compact, interpretable TP prediction framework (172K parameters, 0.34M FLOPs, 2.16 MB) that can be deployable on edge-AI platforms such as wearables and on-board units for real-time monitoring. By operationalizing SSA and kinetic energy principles in ITS interventions, MTPS bridges methodological innovation and policy relevance, advancing the design of cognitively adaptive and harm-aware mobility systems. Future work will validate MTPS in naturalistic riding, extend to multimodal signals, and test adaptive thresholding for real-world ITS interventions.
VIII-A Future Work
Future research will validate MTPS on large-scale, naturalistic data across diverse populations and road environments. Multimodal signals (physiological, behavioral, contextual) will be integrated to strengthen robustness, while adaptive ITS interventions (haptics, throttle modulation, context-aware alerts) will be evaluated for their ability to continuously regulate both crash risk and kinetic energy transfer. Transfer learning and domain adaptation will support scalability across vehicle types, regions, and cultures. Ultimately, MTPS provides a pathway to operationalize the SSA principles (accepting that human error is inevitable but preventable harm is not) by embedding proactive cognitive load management into the design of next-generation ITS.
@@ Acknowledgments
Sumit S. Shevtekar: Conceptualization, Methodology, Data curation, Formal analysis, Software, Visualization, Writing – original draft.Chandresh K. Maurya: Supervision, Methodology, Conceptualization, Validation, Data curation, Writing – review & editing. Gourab Sil: Supervision, Methodology, Conceptualization, Data curation, Writing – review & editing. Subasish Das: Writing – review & editing.
@@
The dataset generated and analyzed during this study is available from the corresponding author upon reasonable request.

### v3

@@ VII-B Application in ITS Interventions
The predicted TP probabilities are thresholded to classify riders into discrete states (NTP, LTP, HTP), as shown in Table XII and Fig. 8. This state-based risk index reframes ITS from passive monitoring to proactive, preventive interventions. Under the SSA, the emphasis is not only on preventing crashes but also on minimizing harm when crashes occur. By integrating MTPS-predicted TP states into adaptive human–machine interfaces (haptics, audio, AR) and vehicle-to-infrastructure (V2I) signaling (e.g., road side unit-based speed guidance, dynamic signal phasing, or targeted alerts), interventions can be triggered before risky behavior escalates. This enables systems to modulate kinetic energy exposure in real time (e.g., slowing riders earlier, guiding safer gap acceptance, or synchronizing vehicle–infrastructure interactions) thereby reducing the severity of crashes even when they are unavoidable. MTPS thus translates cognitive stress estimation into actionable intelligence for SSA-compliant ITS, where thresholds are aligned with both crash likelihood and energy transfer risk, supporting holistic harm reduction strategies.
@@ VIII Discussions and Conclusions
Building on these empirical insights that demonstrate the systemic impact of TP on riding safety, we developed MTPS, a DL architecture that achieves 91.53% accuracy and 98.93% ROC AUC in classifying TP states, outperforming six baselines with only 172K parameters, 0.66 MB model size, and 0.21 ms inference on CPU. Beyond classification accuracy, MTPS demonstrates significant practical utility by improving downstream crash-risk prediction. Using MTPS-predicted TP as a feature improves collision risk accuracy for both Informer (91.25% to 93.51%) and TimesNet (92.10% to 93.90%), approaching oracle performance (93.72% and 94.06%, respectively). Additional validation confirms perfect HTP discrimination (AUC = 1.00) and reliable confidence calibration essential for real-world deployment. MTPS operates on sensor data from individual riders, making it adaptive to rider-specific behavior without assuming homogeneity. Additionally, TP manifests through universal kinematic signals—increased speed, sudden braking, and rapid acceleration—that are consistent across driving environments.
In conclusion, MTPS is a compact, interpretable TP prediction framework (172K parameters, 0.34M FLOPs, 0.66 MB) that can be deployed on edge-AI platforms such as wearables and on-board units for real-time monitoring. By operationalizing SSA and kinetic energy principles in ITS interventions, MTPS bridges methodological innovation and policy relevance, advancing the design of cognitively adaptive and harm-aware mobility systems. Future work will validate MTPS in naturalistic riding, extend to multimodal signals, and test adaptive thresholding for real-world ITS interventions.
VIII-A Limitations and Future Work
This study has several limitations. First, our sample includes only male participants, reflecting Indian demographic patterns where males account for 85.2–87.3% of PTW fatalities, but this limits the generalizability to female riders, as TP perception and riding behavior may differ by gender. Second, data were collected in a controlled simulator environment, which may not fully capture real-world dynamics such as fatigue, weather, or social pressures. Collecting real-world TP data is challenging due to safety and ethical constraints. To address this, we plan to collaborate with commercial riding platforms to collect naturalistic riding data and validate MTPS in real traffic scenarios.
Future research will validate MTPS on large-scale simulator and naturalistic data across diverse populations and road environments. Multimodal signals (physiological, behavioral, contextual) will be integrated to strengthen robustness, while adaptive ITS interventions (haptics, throttle modulation, context-aware alerts) will be evaluated for their ability to continuously regulate both crash risk and kinetic energy transfer. Transfer learning and domain adaptation will support scalability across vehicle types, regions, and cultures. Ultimately, MTPS provides a pathway to operationalize the SSA principles (accepting that human error is inevitable but preventable harm is not) by embedding proactive cognitive load management into the design of next-generation ITS.
@@ Acknowledgments
Sumit S. Shevtekar: Conceptualization, Methodology, Data curation, Formal analysis, Software, Visualization, Writing – original draft. Chandresh K. Maurya: Supervision, Methodology, Conceptualization, Validation, Data curation, Writing – review & editing. Gourab Sil: Supervision, Methodology, Conceptualization, Data curation, Writing – review & editing.
@@
All data and code used will be made available from the corresponding author on a reasonable request upon publication.